OpenStudy (anonymous):

I will give medal!!! Which expression does not result in a value of eighty-eight point four one? fifty-three point nine plus thirty-four point five one ninety point six three minus two point two two three hundred point four minus two hundred eleven point nine nine seventy point one plus eighteen point four

OpenStudy (timmyk14):

Seriously dude

OpenStudy (anonymous):

could you write that out please

OpenStudy (anonymous):

I'm giving out medals

OpenStudy (anonymous):

88.41

OpenStudy (timmyk14):

It's B. 90.63-2.2= 88.43

OpenStudy (timmyk14):

Oops. didn't see it said 2.22

OpenStudy (timmyk14):

It's not B xD

OpenStudy (anonymous):

53.9+34.51= 90.63+2.22= 300.4-211.9= 70.1+18.4=

OpenStudy (timmyk14):

It's the last one.

OpenStudy (anonymous):

yeah, that's right

OpenStudy (anonymous):

a. 88.41

OpenStudy (anonymous):

Well, you need to solve each individual one cx

OpenStudy (anonymous):

b.92.85

OpenStudy (anonymous):

c.88.5

OpenStudy (timmyk14):

90.6-2.22=92.85????

OpenStudy (anonymous):

88.4

OpenStudy (anonymous):

oh sorry lol i thought it was plus

OpenStudy (anonymous):

88.38

OpenStudy (anonymous):

so yea its 88.41

OpenStudy (anonymous):

Lol typed up wrong but his question is which isn't

OpenStudy (anonymous):

it is the ladt one

OpenStudy (anonymous):

last*

OpenStudy (anonymous):

no because 300.4-211.9 is 88.5 as well

OpenStudy (timmyk14):

Holy sh*t guys... A- 53.9 + 34.51 = 88.41 B- 90.63 - 2.22 = 88.41 C- 300.4 - 211.99 = 88.41 D- 70.1 + 18.4 = 88.5 ****ANSWER IDIOTS***

OpenStudy (anonymous):

Oh, its .99

OpenStudy (timmyk14):

You have me facepalming the sh*t out of my head right now dude